Monday 04/09/12

10:00 A – 10:30 P

  • Lab meeting (Doory)
  • Fix alignment of 750 laser with Colenso’s help
  • Freeze Bcd-GFP and eve-GFP bottles.
  • NuGru presentation (Andi)
  • Working on snail manuscript supplemental figure captions
  • Download ModEncode data for H3K4me3 and H3K27ac
  • Etch and mount new slide of Psc K4 stained embryos for STORM (6:00P)
  • O/N STORM imaging of Psc K4 (after 2 calibration runs.  Started at 9:00P).
  • O/N PCR to clone Hox enhancers and Hox introns (13 reactions)
  • Retransform DH10alpha cells with Taf and Tub ligations (14 uL of cells + 5 uL of ligation reaction).  Recovered in 200 uL SOC for 1 hr.
  • Plated 60 uL of cells on Amp plates with  60 uL IPTG 60 uL XGAL preabsorbed.  Spread with glass beads.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.